What is the molar mass of a gas, if 12.2g of the gas occupies a volume of 4.5L at 1.76atm and 22 degrees celsius?
Answer #1

PV = nRT

where, P = pressure = 1.76 atm

V = volume = 4.5 L

n = number of moles

R = Gas constant

T = temperature = 22 + 273 = 295 K

1.76 * 4.5 = n * 0.0821 * 295

7.92 = n * 24.2

n = 7.92 / 24.2 = 0.327 mole

number of moles = mass of gas / molar mass of gas

0.327 mole = 12.2 g / molar mass of gas

molar mass of gas = 12.2 g / 0.327 mole = 37.3 g/mol

Therefore, the molar mass of gas = 37.3 g/mol